OpenStudy (anonymous):

Assume that there are a series of emissions for the hydrogen atom in which n(subscript f) = 2, find the electromagnetic spectrum in which these emissions are observed.

OpenStudy (aaronq):

use the rydberg equation to find the wavelengths of light correspondent to transitions from n>2 to n=2. \(\dfrac{1}{\lambda}=R \;(\dfrac {1}{n_f^2}-\dfrac{1}{n_i^2})\ \;\;;\;\;\;\;\; where,\;R=\dfrac{m_ee^4}{(4\pi)^3\epsilon_0^2\hbar^3c }\approx1.097*10^7 m^{-1}\)

OpenStudy (anonymous):

Hi, thanks for the answer, what should I plug in for n1 and n2?

OpenStudy (anonymous):

Hi, thanks for the answer, what should I plug in for n1 and n2?

OpenStudy (aaronq):

Emission implies that \(n_i>n_f\), and in the question it's telling you that \(n_f\)=2, so \(n_i\)>2.

OpenStudy (aaronq):

the question is asking for a spectrum; in other words, a range.